OpenStudy (anonymous):

-7/8 divided by - 3/80

OpenStudy (anonymous):

\[\frac{-\frac{7}{8}}{-\frac{3}{80}}\]

OpenStudy (anonymous):

should be 70/3

OpenStudy (anonymous):

yes 70/3

OpenStudy (chriss):

(-7/8)*(-80/3)=560/24=70/3